# De-Hallucinator: Iterative Grounding for LLM-Based Code Completion

### v1

@@ Abstract.
Large languages models (LLMs) trained on datasets of publicly available source code have established a new state-of-the-art in code completion. However, these models are mostly unaware of the code that already exists within a specific project, preventing the models from making good use of existing APIs. Instead, LLMs often invent, or “hallucinate”, non-existent APIs or produce variants of already existing code. Although the API information is available to IDEs, the input size limit of LLMs prevents code completion techniques from including all relevant context into the prompt. This paper presents De-Hallucinator, an LLM-based code completion technique that grounds the predictions of a model through a novel combination of retrieving suitable API references and iteratively querying the model with increasingly suitable context information in the prompt. The approach exploits the observation that LLMs often predict code that resembles the desired completion, but that fails to correctly refer to already existing APIs. De-Hallucinator automatically identifies project-specific API references related to the code prefix and to the model’s initial predictions and adds these references into the prompt. Our evaluation applies the approach to the task of predicting API usages in open-source Python projects. We show that De-Hallucinator consistently improves the predicted code across four state-of-the-art LLMs compared to querying the model only with the code before the cursor. In particular, the approach improves the edit distance of the predicted code by 23–51% and the recall of correctly predicted API usages by 24–61% relative to the baseline.
copyright: none
@@ 1. Introduction
Large language models (LLMs) have proven effective in many natural language (Brown et al., 2020) and programming tasks (Chen et al., 2021; Bareiß et al., 2022; Xia and Zhang, 2022; Poesia et al., 2022; Jain et al., 2022; Xu et al., 2022). One of the most promising tasks is code completion, as evidenced by the rapid adoption of LLM-based code completion tools, such as Copilot [^1] https://github.com/features/copilot and Tabnine [^2] https://www.tabnine.com/. Since these models are trained on large corpora of code, they usually generate syntactically correct code, and sometimes even the exact code that the developer intends to write.
A crucial factor for obtaining the desired completions is the prompt, i.e., the input given to the LLM. State-of-the-art LLMs build on transformers (Vaswani et al., 2017), which use self-attention to generate sequences of tokens in an auto-regressive process. That is, the model decides what token to predict next based on the tokens in the prompt and any already generated tokens. Hence, designing effective prompts, sometimes called prompt engineering, is a crucial part of developing a practical LLM-based technique (Liu et al., 2021; Shrivastava et al., 2023; Nashid et al., 2023).
For code completion, the prompt typically consists of the incomplete code, and optionally, some additional context. The incomplete code can be at any level of granularity, such as a partial line, an incomplete function, or an incomplete file. The context can be any textual information that may help the model to predict a suitable completion (Jain et al., 2022; Schäfer et al., 2023). For example, a commonly used prompt is a fixed-size window of code just before the cursor (Chen et al., 2021; Xu et al., 2022), i.e., code that a developer has already written and that may contain hints about what variables, functions, etc. to use in the completion.
Despite the impressive success of LLM-based code completion, these techniques are still at an early stage. In particular, we identify two key challenges faced by current approaches.
Challenge 1: Project-specific APIs. As LLMs are trained on huge code bases, they effectively capture typical language idioms and commonly used libraries. In contrast, a general-purpose model lacks knowledge of project-specific APIs, and may fail to correctly use existing functions and classes. In particular, this lack of knowledge may cause the model to “hallucinate” APIs that actually do not exist in the current code base (Nguyen and Nadi, 2022), or perhaps even worse, it may reimplement some functionality that is already present in the code base.
@@
class DataStore():
"""Data structure for storing documents."""
def __init__(self, file: str):
with open(file, r’) as f:
self.documents = f.read().split(’—–’)
def find_by_keyword(self, keyword: str) -> List[str]:
"""Returns all documents that contain the keyword."""
return [d for d in self.documents if keyword in d]
def relevance(document: str, keyword: str) -> float:
"""Returns the relevance of the document to the keyword."""
return document.count(keyword) / len(document)
def search(ds: DataStore, keyword: str, top_k: int) -> List[str]:
%**** main.tex Line 125 ****"""Returns the top_k most relevant documents that contain the keyword sorted by relevance."""
docs = ds.find_by_keyword(keyword)
[…]
@@
def search(ds: DataStore, keyword: str, top_k: int) -> List[str]:
"""Returns the top_k most relevant documents that contain the keyword sorted by relevance."""
docs = ds.find_by_keyword(keyword)
[…]
@@
def search(ds: DataStore, keyword: str, top_k: int) -> List[str]:
%**** main.tex Line 150 ****"""Returns the top_k most relevant documents that contain the keyword sorted by relevance."""
docs = ds.find_by_keyword(keyword)
[…]
Figure 3. The completion of search by ChatGPT highlighted in gray, and the wrong API usage highlighted in red.
@@
Requesting an LLM, e.g., CodeGen (Nijkamp et al., 2022), to complete the search function given a prompt that contains all existing code in UI.py results in Fig. 2. The code is partially correct, but refers to a non-existing API (an attribute doc.relevance). A larger model, e.g., ChatGPT, faces similar challenges and implements search as shown in Fig. 3. As before, the code is partially correct but invokes an API not available in the project (a function compute_relevance_score). The underlying problem for both models is that they are not aware of the project-specific APIs that should be used to complete the code, and hence, the LLMs simply hallucinate some plausible but ultimately wrong APIs.
@@
Challenge 2: Prioritizing context. A naive solution to address Challenge 1 would be to simply add all of the code in the project into the prompt. However, LLMs have a fixed maximum sequence length, which restricts how many tokens one can provide to the model. Choosing the most helpful context for a given completion task is crucial, but an inherently difficult problem, because the optimal context depends on the desired code, which is not known a-priori. While traditional code completion approaches typically have access to various kinds of information available in IDEs, such as the names and types of program elements, providing all this information, or even all the code of the project, to an LLM is impossible due to the limited prompt size.
@@
This paper presents De-Hallucinator, which addresses the above challenges through a novel combination of retrieval-augmented prompts and an iterative form of LLM-based code completion. Our approach uses three types of prompts, which provide increasingly suitable context information. At first, De-Hallucinator queries the LLM with the conventional prompt that contains only the preceding code. Because the preceding code alone often is insufficient to obtain a correct completion, the approach then augments the prompt with API references that are most similar to the preceding code, which makes the second prompt type. However, also this prompt may fail to find a suitable completion, because the preceding code might not be similar to the desired API, or there are other APIs more similar to the preceding code than the correct one. Since LLMs often predict code that resembles the desired completion, but fail to correctly refer to project-specific APIs, the completion from the previous prompt types often resembles the desired API. Hence, to construct the third type of prompt, De-Hallucinator leverages these hints about what code the model intends to predict to retrieve suitable project-specific APIs, which are then added to the third type of prompt.
The result of these queries to the LLMs is a sequence of completions likely to be increasingly suitable. In a practical deployment, these completions could be shown to a user as a ranked list, or be further filtered by a validation mechanism, such as running a test suite.
Our approach addresses the two challenges described above. To address Challenge 1, De-Hallucinator adds information about project-specific APIs into the prompt. As a result, the model avoids (re)inventing code and instead correctly invokes existing APIs. The idea of augmenting an LLM with well-grounded facts relates to work on grounding of language models for natural languages (Roy, 2005; Ahn et al., 2022; Gu et al., 2022). To the best of our knowledge, we are the first to apply grounding to the problem of predicting API usages in code. To address Challenge 2, De-Hallucinator uses the model itself to gather hints about what kind of context information is most suitable to help the model make a better prediction. This iterative approach complements prior work that tries to guess the most suitable context from the incomplete code alone (Shrivastava et al., 2023; Ding et al., 2022). Our work also relates to other retrieval-based prompt augmentation techniques, which try to find suitable few-shot examples (Nashid et al., 2023) . In contrast to existing retrieval-based prompt augmentation, De-Hallucinator retrieves API references instead of few-shot examples and addresses the task of code completion.
@@
The presented approach offers several benefits. First, De-Hallucinator works with any off-the-shelf LLM trained on code, because the approach treats the model as a black box. In particular, we do not require to train or fine-tune the model in any way, but simply exploit the fact that its predictions contain implicit hints about additional context the model would benefit from. Second, because APIs usually evolve only slowly, De-Hallucinator can pre-compute, and occasionally update in the background, the set of project-specific API references. As a result, the latency of code completion is not impacted by any expensive program analysis, which is important for practical adoption. Finally, the approach is fully transparent to developers, because the approach hides the iterative interaction with the LLM from the user and simply returns a ranked list of predictions.
@@
We evaluate De-Hallucinator by applying the approach to four state-of-the-art LLMs for code, namely CodeGen (Nijkamp et al., 2022), CodeGen 2.5 (Nijkamp et al., 2023), UniXcoder (Guo et al., 2022a), and StarCoder+ (Li et al., 2023a). Although, conceptually, the approach can be applied to any programming language, our evaluation focuses on Python as it is one of the most popular languages  [^3] https://octoverse.github.com/2022/top-programming-languages and a common target of prior work on code completion (Li et al., 2018; Svyatkovskiy et al., 2019; Chen et al., 2021; Guo et al., 2022a; Nijkamp et al., 2022, 2023; Li et al., 2023a; Zhang et al., 2023b). Compared to querying the model with a prompt that contains only the code before the missing lines, we find that De-Hallucinator enables the model to provide more accurate predictions. In particular, we show a relative improvement of 23.28-50.64% in edit distance, of 12.12-27.48% in normalized edit similarity, and of 23.90-60.98% in recall of correctly predicted API usages.
@@
  • Empirical motivation showing that the problem of predicting API usages affects a large portion of failed code completion tasks.
@@
  • A novel algorithm that combines retrieval-augmented prompting with an iterative method for constructing increasingly suitable prompts by using the hallucinations produced in earlier iterations to augment the context information provided in the prompts of future iterations.
  • Empirical evidence that, across four state-of-the-art LLMs, De-Hallucinator offers more accurate completions than querying the same models with a fixed prompt.
@@ 2.1. Project-Specific APIs
The main motivation for this work is our observation that LLMs often hallucinates code that resembles the desired completion, but that fails to correctly refer to an API. To assess the importance of this limitation, we manually investigate and classify the reasons why an LLM fails to predict the desired completion. We perform this preliminary study on 50 function-level code completion tasks, which we collect by (i) randomly selecting ten Python projects from a curated list of open-source projects  [^4] https://github.com/vinta/awesome-python. We randomly sample ten application domains and then sample one project from each domain. and (ii) by then randomly selecting five functions from each project. The only filtering we perform is to ignore functions with more than 25 lines, as these are likely out of reach for today’s LLMs. For each of the 50 functions, we query an LLM (CodeGen 2.5 with 7B parameters and 4-bit quantization) with the code before the beginning of the function body, including the function signature and any docstring, in the prompt.
Given the 50 pairs of an LLM-predicted function body and the ground-truth function body, we manually classify them based on two questions. First, is the prediction correct w.r.t. the ground truth, where “correct” includes exact matches and semantically equivalent code? Second, does the ground truth contain an API usage, e.g., a function call, that is missing in the prediction? Initially, two of the authors independently classify the 50 pairs, with an inter-rater agreement (Cohen’s kappa) of 0.76, which is considered excellent (Fleiss, 1981). The two authors then discuss the few cases with initial disagreement. The initial disagreement in each of the cases proved to be due to a human error, such as missing the difference between similar identifier names, high and highpass, or similar literal values, 3 and 8. After resolving these errors, the authors reach a consensus about all 50 pairs.
The final inspection results show that in 13 out of the 50 cases, the LLM either predicts exactly the expected function body or a function body that is semantically equivalent to the expected one. For 22 out of the 37 remaining cases, there is at least one API usage that the LLM fails to correctly predict, similar to the examples in Fig. 2 and Fig. 3. In other words, the problem identified and addressed in this work affects 44% of all studied function-level code completion tasks, and even 59% of all tasks where the LLM alone fails to predict the expected code.
@@ 2.2. Prioritizing Context
To validate the importance of the second challenge, we compare the amount of code in a single project to the prompt sizes of high-end LLMs. The models in the popular GPT series by OpenAI have prompt sizes between 2,048 (GPT-3 models) and 32,768 (GPT-4-32k) tokens. In contrast, in a sample dataset of 50 Python projects, which are randomly selected from the same curated list of projects as above, there are 488,635 tokens per project, on average. Furthermore, the average project has around 13 files longer than 8,192 tokens, and 22 projects in our sample have at least one file longer than 32,768 tokens. This means that even knowing the exact file that contains the relevant context (e.g., based on heuristics, such as recently used files or similar file names) leaves us with more tokens than one could fit into the prompt. In other words, simply adding all potentially relevant code to the prompt is not a viable solution, but we need to prioritize the context information.
@@ 3. Approach
This section describes an LLM-based code completion approach that uses retrieval of relevant APIs to improve the prompt for additional queries to the model. We call the approach De-Hallucinator, as it reduces the hallucinations of the LLM by providing relevant API references to ground the model. After defining the problem that De-Hallucinator addresses (Section 3.1), we provide an overview of the approach and present its main algorithm (Section 3.2). Sections 3.3 to 3.6 then present each of the components of De-Hallucinator in detail.
3.1. Problem Definition
Before describing our approach, we define the problem that De-Hallucinator addresses. Given an incomplete piece of code $c$, where some code is missing at a cursor location $l$, the problem is to predict code $c^{\prime}$ to be placed into $c$ at location $l$. Specifically, we consider the problem where $c^{\prime}$ consists of one or more missing lines that contain an API usage, and where $c$ is all the code in the project except for the code between $l$ and the end of the current function.
This problem definition matches the common scenario of a developer implementing a function in an existing project, where the code to be written should use a project-specific API. Note that the API usage does not have to start directly at the cursor location, but only to start in the next line. For example, suppose that the cursor in Fig. 1 is at the beginning of the code marked with gray background. Everything above the cursor is our incomplete code $c$, and the problem is to predict the marked code $c^{\prime}$, which refers, e.g, to the project-specific relevance API.
To address this problem, we assume to have a generative language model $m$ trained on a large corpus of code. The code $c$ is part of a project $p$, which may contain a lot of already written code that $c^{\prime}$ could refer to. However, we assume that the existing code in $p$ largely exceeds the maximum input size accepted by $m$. We further assume that the existing code in $p$ has not been part of $m$’s training data, which matches the practical usage scenario of using a single large-scale model as the basis of a code completion tool used by many projects.
3.2. Overview and Main Algorithm

### v3

# frame=lines,breaklines,breaksymbol= ,fontsize=,escapeinside=||,highlightcolor=bggray

@@ Abstract.
Large language models (LLMs) trained on datasets of publicly available source code have established a new state of the art in code generation tasks. However, these models are mostly unaware of the code that exists within a specific project, preventing the models from making good use of existing APIs. Instead, LLMs often invent, or “hallucinate”, non-existent APIs or produce variants of already existing code. This paper presents De-Hallucinator, a technique that grounds the predictions of an LLM through a novel combination of retrieving suitable API references and iteratively querying the model with increasingly suitable context information in the prompt. The approach exploits the observation that predictions by LLMs often resemble the desired code, but they fail to correctly refer to already existing APIs. De-Hallucinator automatically identifies project-specific API references related to the model’s initial predictions and adds these references into the prompt. Unlike retrieval-augmented generation (RAG), our approach uses the initial prediction(s) by the model to iteratively retrieve increasingly suitable API references. Our evaluation applies the approach to two tasks: predicting API usages in Python and generating tests in JavaScript. We show that De-Hallucinator consistently improves the generated code across five LLMs. In particular, the approach improves the edit distance by 23.3–50.6% and the recall of correctly predicted API usages by 23.9–61.0% for code completion, and improves the number of fixed tests that initially failed because of hallucinations by 63.2%, resulting in a 15.5% increase in statement coverage for test generation.
@@ 1. Introduction
Large language models (LLMs) have proven effective in many natural language (Brown et al., 2020) and programming tasks (Chen et al., 2021; Bareiß et al., 2022; Xia and Zhang, 2022; Poesia et al., 2022; Jain et al., 2022; Xu et al., 2022; Schäfer et al., 2024). Rapid adoption of LLM-based tools, such as Copilot [^1] https://github.com/features/copilot and Tabnine [^2] https://www.tabnine.com/, shows practical productivity benefits (Ziegler et al., 2024; Liang et al., 2024). State-of-the-art LLMs build on transformers (Vaswani et al., 2017), which use self-attention to generate sequences of tokens in an auto-regressive process. That is, the model decides what token to predict next based on the tokens in the prompt and any already generated tokens. Hence, designing effective prompts, sometimes called prompt engineering, is a crucial part of developing a practical LLM-based technique (Liu et al., 2021; Shrivastava et al., 2023; Nashid et al., 2023).
Despite the impressive success of LLM-based code generation, these techniques are still at an early stage. In particular, we identify two key challenges faced by current approaches:
Challenge 1: Project-specific APIs. As LLMs are trained on huge code bases, they effectively capture typical language idioms and commonly used libraries. In contrast, a general-purpose model lacks knowledge of project-specific APIs, and may fail to correctly use existing functions and classes. In particular, this lack of knowledge may cause the model to “hallucinate” APIs that actually do not exist in the current code base (Nguyen and Nadi, 2022), or perhaps even worse, it may reimplement some functionality that is already present in the code base. Developers perceive this lack of knowledge about project-specific APIs as an obstacle to using AI programming assistants (Liang et al., 2024).
@@
{minted}
[label=DataStore.py]python class DataStore(): def __init__(self, file: str): with open(file, ’r’) as f: self.documents = f.read().split(’—–’) … def find_by_keyword(self, keyword: str) -> List[str]: return [d for d in self.documents if keyword in d] … {minted}[label=utils.py]python … def relevance(document: str, keyword: str) -> float: return document.count(keyword) / len(document) … {minted}[label=UI.py,highlightlines=4]python … def search(ds: DataStore, keyword: str, top_k: int) -> List[str]: docs = ds.find_by_keyword(keyword) return sorted(docs, key=lambda d: relevance(d, keyword), reverse=True)[:top_k] …
@@
{minted}
[highlightlines=3]python def search(ds: DataStore, keyword: str, top_k: int) -> List[str]: docs = ds.find_by_keyword(keyword) return sorted(docs, key=lambda x: |x.score,|, reverse=True)[:top_k]
@@
Requesting an LLM, e.g., CodeGen (Nijkamp et al., 2022), to complete the search function given a prompt that contains all existing code in UI.py results in Fig. 2. The code is partially correct, but refers to a non-existing API (an attribute x.score). The underlying problem is that the models are not aware of the project-specific APIs that should be used to complete the code, and hence, the LLM simply hallucinates some plausible but ultimately wrong APIs.
@@
Challenge 2: Prioritizing context. A naive solution to address Challenge 1 would be to simply add all of the code in the project into the prompt. However, LLMs have a fixed maximum sequence length, which restricts how many tokens one can provide to the model. Even with the recent increases in sequence length of LLMs, providing the most useful context can improve the output and reduce the costs. Choosing the most helpful context for a given completion task is crucial, but an inherently difficult problem, because the optimal context depends on the desired code, which is not known a-priori. While traditional code completion approaches typically have access to various kinds of information available in IDEs, such as the names and types of program elements, providing all this information, or even all the code of the project, to an LLM is impossible due to the limited prompt size.
@@
This paper presents De-Hallucinator, which addresses the above challenges through a novel combination of retrieval-augmented generation (RAG) and an iterative form of LLM-based code generation. Our approach uses three types of prompts, which provide increasingly suitable context information. The initial prompt type is querying the LLM with the conventional prompt, i.e., without any retrieval. Retrieval-augmented generation (RAG) (Lewis et al., 2020) proposes to retrieve relevant context based on the initial prompt to address both Challenges 1 and 2, which we refer to as the RAG prompt type. The idea of augmenting an LLM with well-grounded facts relates to work on grounding of language models for natural languages (Roy, 2005; Ahn et al., 2022; Gu et al., 2022). However, also this prompt may fail to generate factually correct code (i.e., without hallucinations), because the initial prompt might not have any similar code to the desired API, or there are other APIs more similar to the initial prompt than the correct one. We make the improtant observation that the generated code from the previous prompt types often resembles the desired API. Hence, we construct a new type of prompt called the iterative prompt. De-Hallucinator leverages the code that the model predicts to retrieve suitable project-specific APIs, which are then added to the iterative prompt for the next round of model prediction. The iterative prompt type complements prior work that tries to guess the most suitable context from the incomplete code alone (Shrivastava et al., 2023; Ding et al., 2022).
@@
The presented approach offers several benefits. First, De-Hallucinator works with any off-the-shelf LLM trained on code because the approach treats the model as a black box. In particular, we do not require to train or fine-tune the model in any way, but simply exploit the fact that its predictions contain implicit hints about additional context the model would benefit from. Second, because APIs usually evolve only slowly, De-Hallucinator can pre-compute, and occasionally update in the background, the set of project-specific API references. As a result, the latency of code generation is not impacted by any expensive program analysis, which is important for practical adoption. Finally, the approach is fully transparent to developers, because the approach hides the iterative interaction with the LLM from the user and simply returns a ranked list of predictions.
@@
We evaluate De-Hallucinator by applying the approach to code completion with four state-of-the-art LLMs for code, namely CodeGen (Nijkamp et al., 2022), CodeGen 2.5 (Nijkamp et al., 2023), UniXcoder (Guo et al., 2022a), and StarCoder+ (Li et al., 2023a), and to test generation with GPT-3.5-turbo. Conceptually, the approach can be applied to any programming language, and our evaluation focuses on two popular languages, Python and JavaScript as they are among the most popular languages  [^3] https://octoverse.github.com/2022/top-programming-languages and common targets of prior work on code completion (Li et al., 2018; Svyatkovskiy et al., 2019; Chen et al., 2021; Guo et al., 2022a; Nijkamp et al., 2022, 2023; Li et al., 2023a; Zhang et al., 2023b) and test generation (Arteca et al., 2022; Schäfer et al., 2024). Compared to conventional prompts, we find that De-Hallucinator enables the models to provide more accurate predictions. In particular, we show a relative improvement of 23.3–50.6% in edit distance, and of 23.9–61.0% in recall of correctly predicted API usages for code completion. Moreover, we show relative improvement of 17.9% in number of passing tests, of 15.5% in coverage, and of 63.2% in the number of mitigated hallucinations.
@@
  • Empirical motivation showing that API hallucinations affect a large portion of failed code completion and test generation tasks.
@@
  • A novel algorithm that combines retrieval-augmented generation with an iterative method for constructing increasingly suitable prompts by using the hallucinations produced in earlier iterations to augment the context information provided in the prompts of future iterations.
  • Empirical evidence that, across two code generation tasks, two programming languages, and five state-of-the-art LLMs, De-Hallucinator offers more accurate generations than conventional prompts.
@@ 2.1. Project-Specific APIs
The main motivation for this work is our observation that LLMs often hallucinate code that resembles the desired code, but that fail to correctly refer to an API. To assess the importance of this limitation, we investigate the prevalence and causes of hallucinated APIs in the two code generation tasks focused in this paper. For code completion, we manually investigate and classify the reasons why an LLM fails to predict the desired completion. We perform this preliminary study on 50 function-level code completion tasks, which we collect by (i) randomly selecting ten Python projects from a curated list of open-source projects  [^4] https://github.com/vinta/awesome-python. We randomly sample ten application domains and then sample one project from each domain. and (ii) by then randomly selecting five functions from each project. The only filtering we perform is to ignore functions with more than 25 lines, as these are likely out of reach for today’s LLMs. For each of the 50 functions, we query an LLM (CodeGen 2.5 with 7B parameters and 4-bit quantization) with the code before the beginning of the function body, including the function signature and any docstring, in the prompt.
Given the 50 pairs of an LLM-predicted function body and the ground-truth function body, we manually classify them based on two questions. First, is the prediction correct w.r.t. the ground truth, where “correct” includes exact matches and semantically equivalent code? Second, does the ground truth contain an API usage, e.g., a function call, that is missing in the prediction? Initially, two of the authors independently classify the 50 pairs, with an inter-rater agreement (Cohen’s kappa) of 0.76, which is considered excellent (Fleiss, 1981), and after discussing the discrepencies reach a consensus about all 50 pairs.
The final inspection results show that in 13 out of the 50 cases, the LLM either predicts exactly the expected function body or a function body that is semantically equivalent to the expected one. For 22 out of the 37 remaining cases, there is at least one API usage that the LLM fails to correctly predict, similar to the example in Fig. 2. In other words, the problem identified and addressed in this work affects 44% of all studied function-level code completion tasks, and even 59% of all tasks where the LLM alone fails to predict the expected code.
For test generation, we use error messages of crashing tests generated by TestPilot (Schäfer et al., 2024) on a diverse set of 12 JavaScript projects. We automatically count the number of generated tests that result in “* is not a function”, or “Cannot read properties of undefined” errors, which typically indicates hallucinations of non-existing APIs. We find that, on average, 16.4% of all generated tests fail because of the aforementioned errors, which indicates that hallucinations of non-existing APIs are a common problem in test generation as well.
@@ 2.2. Prioritizing Context
To validate the importance of the second challenge, we compare the amount of code in a single project to the prompt sizes of high-end LLMs. The models in the popular GPT series by OpenAI have prompt sizes between 4k (GPT-3.5 models) and 128k (GPT-4) tokens. In contrast, in a sample dataset of 50 Python projects, which are randomly selected from the same curated list of projects as above, there are 488,635 tokens per project, on average. Furthermore, the average project has around 13 files longer than 8,192 tokens, and 22 projects in our sample have at least one file longer than 32,768 tokens. This means that even knowing the exact file that contains the relevant context (e.g., based on heuristics, such as recently used files or similar file names) leaves us with more tokens than one could fit into the prompt of some models. Even for models with longer context window, considering a cost of 0.5\$ for 100k tokens means that the cost of long prompts would be impractical for regular use. In other words, simply adding all potentially relevant code to the prompt is not a viable solution, but we need to prioritize the context information.
@@ 3. Approach
This section describes our approach for iteratively retrieving relevant APIs to improve the prompts for code generation tasks. We call the approach De-Hallucinator, as it reduces the hallucinations of the LLM by providing relevant API references to ground the model. First, we provide an overview of the approach (Section 3.1), and then present each of the components of De-Hallucinator in detail (Sections 3.2 to 3.5).
3.1. Overview